OpenStudy (anonymous):

Can someone help me in Algebra 2? Need you to help me with a few problems! Let me know :)

OpenStudy (gorv):

???????????????/

OpenStudy (anonymous):

Solve the absolute value equation /3x+2/ = 2

OpenStudy (anonymous):

a. x=1 or x= 4/3 b. x=0 or x=4/3 c. x=0 or x= -4/3 d. x=1 or x = -4/3

OpenStudy (gorv):

|3x+2|=2 on removing mode 3x+2=2 or 3x+2=-2 3x=0 3x=-4 x=0 x=-4/3

OpenStudy (gorv):

@kkaylaburton

OpenStudy (anonymous):

Thank you! okay now, What is an equation of the line, in point-slope form, that passes through the given point and has the given slope? point; (7,3); slope: 4/7

OpenStudy (anonymous):

a. y - 3= 4/7 (x-7) b. y - 7 = 4/7 (x-3) c. y - 3 = 4/7 (x+7) d. y - 3 = -4/7 (x-7)

OpenStudy (gorv):

y-3=4/7 (x-7)

OpenStudy (anonymous):

you research the cost of a gallon of gasoline over several years to look for a trend. the table shows yours data. the line of best fit uses years since 1980 as input value. what is a line of best ft for the data? based upon these prices how much would you expect to pay in 2028?

OpenStudy (anonymous):

year: 1980 1984 1988 1992 1996 200 2004 2008 2012 price per gallon: $1.25 $1.27 $1.05 $1.28 $1.43 $1.68 $2.11 $3.72 $3.99

OpenStudy (anonymous):

a. y= 0.087x + 2.26; $6.44 b. y= 0.087x + 0.587; $4.76 c. y= 0.134x + 1.25; $7.68 d. y= 0.0347x + 0.856; $2.52

OpenStudy (anonymous):

@gorv

OpenStudy (anonymous):

I got B

OpenStudy (anonymous):

okay! thank you! compare the function with the parent function. without graphing, what are the vertex, axis of symmetry, and transformations of the given function? y = /10x-2/ -7

OpenStudy (anonymous):

The vertex would be in (2,-7), The graph moved 2 right and 7 down from the parent function

OpenStudy (anonymous):

Last night, Julie's pet hamster Zoe kept Julie awake for at least an hour running on her exercise wheel and scratching the of her cage. Zoe ran on her wheel at least twice as long as she scratched at the corners of her cage, and she spent more than 1/4 hour running on her wheel. If x is the number of hours Zoe spent running on her wheel and y is the number of hours spent scratching her cage, which of the systems of inequalities represents the time Zoe kept Julie awake last night.